A thin glass rod is bent into a semi circle of radius R a charge Q is uniformly distributed along the upper-half and a charge −Q is uniformly distributed along the lower-half. Find the electric field E~ at the center of the semicircle. 


1
Expert's answer
2021-06-30T10:59:59-0400

The semi circle should look like the one shown below



considering the center P of the semicircle, and taking two elements of width "d\theta" the magnitude of the field at P is;

"dE=\frac{1}{4\pi \epsilon_{o}} \times \frac{rd\theta \times Q(\pi r/2)}{r^2}=\frac{Q}{2\pi ^2\epsilon_{o}r^2} d\theta"

field at x due to pair of electrons"=2dEsin\theta"

"E=_{0}\int^{\pi /2} 2dEsin\theta=2\times _{0}\int^{\pi /2}\frac{Q}{2\pi \epsilon_{o} r^2}sin\theta d\theta=\frac{Q}{\pi ^2 \epsilon_{o}r^2}"